# Proximal Oracles for Optimization and Sampling

### v2

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex but not necessarily smooth.
@@
Optimization and sampling are two of the most important algorithmic tools at the interface of data science and computation. Optimization has been extensively studied in machine learning, communication, and supply chain. Drawing samples from a given (often unnormalized) probability density plays a crucial role in many scientific and engineering problems that face uncertainty (either physically or algorithmically). Sampling algorithms are widely used in many areas such as statistical inference/estimation, operations research, physics, biology, and machine learning, etc [1, 10, 11, 15, 21, 22, 27, 59]. For instance, in Bayesian inference, one draws samples from the posterior distribution to infer its mean, covariance, or other important statistics. Sampling is also heavily used in molecular dynamics to discover new molecular structures.
@@
This work is along the recent line of researches that lie in the interface of sampling and optimization [9, 57]. Indeed, sampling is closely related to optimization. On the one hand, optimization can be viewed as the limit of sampling when the temperature parameter, or equivalently the randomness in the problem, goes to 00. On the other hand, sampling can be viewed as an optimization over the manifold of probability distributions [62, 64]. The popular gradient-based Markov chain Monte Carlo (MCMC) methods such as Langevin Monte Carlo (LMC) [6, 18, 51, 53], Metropolis-adjusted Langevin algorithm (MALA) [2, 52, 53], and Hamiltonian Monte Carlo (HMC) [45] resemble the gradient-based algorithms in optimization and can be viewed as the sampling counterparts of them.
@@
The goal of this paper is to develop efficient proximal algorithms to solve optimization problems (1) as well as to draw samples from potentials (2), where both $f$ in (1) and (2) lack smoothness. In particular, we consider two settings where the convex objective/potential function $f$ is either Hölder smooth (i.e., the (sub)gradient $f^{\prime}$ is Hölder-continuous with exponent $\alpha\in[0,1]$) or a hybrid function with multiple Hölder smooth components. The core of both proximal optimization and sampling algorithms is a proximal map of $f$. Based on an efficient implementation of the proximal map, we develop an adaptive proximal bundle method to solve (1). Building upon the proximal map of $f$ and rejection sampling, we propose a highly efficient method to realize a proximal sampling oracle, which is used in a proximal sampling framework [29, 5] in the same spirit as the proximal point method for optimization. With those proximal oracles for optimization and sampling in hand, we are finally able to establish the complexity to sample from densities with non-smooth potentials.
@@ 2 Proximal optimization and sampling
Starting from the original paper [29] that proposes ASF, a subsequent paper [5] provides an improved analysis of ASF, and [65] also studies Gibbs sampling based on ASF over bipartite graphs. In Algorithm 1, sampling $y_{k}$ given $x_{k}$ in step 1 can be easily done since $\pi^{Y|X}(y\mid x_{k})={\cal N}(x_{k},\eta I)$ is a simple Gaussian distribution. Sampling $x_{k+1}$ given $y_{k}$ in step 2 is however a nontrivial task; it corresponds to the so-called restricted Gaussian oracle (RGO) for $f$ introduced in [29], which is defined as follows.
@@
A rejection sampling-based implementation of RGO for general convex nonsmooth potential function $f$ with bounded Lipschitz constant is given in [35]. If the stepsize $\eta$ is small enough, then it only takes a constant number of rejection steps to generate a sample according to RGO in expectation. Another exact realization of RGO is provided in [36] for nonconvex hybrid potential $f$ satisfying Hölder continuous conditions. It is also shown that the expected number of rejections to implement RGO is a small constant if $\eta$ is small enough. Other inexact realizations of RGO based on approximate rejection sampling are studied in [17, 12]. In all these implementations, a key step is realizing the proximal map (4). It is worth noting that [36] also connects ASF with other well-known Langevin-type sampling algorithms such as Langevin Monte Carlo (LMC) and Proximal Langevin Monte Carlo (PLMC) via RGO. In a nutshell, [36] shows that both LMC and PLC are instances of ASF but with approximate implementations of RGO, which always accept the sample from the proposal distribution without rejection. Hence, this explains why the samples generated by LMC is biased while ASF is unbiased.
@@ 3 Algorithm and complexities for the proximal subproblem
0:  Let $y\in\mathbb{R}^{d}$, $\eta>0$, and $\delta>0$ be given, and set $x_{0}=\tilde{x}_{0}=y$.
  for $j=1,2,\cdots$ do
[…]
$x_{j}$ $=\underset{x\in\mathbb{R}^{d}}{\mathrm{argmin}\,}\left\{f_{j}^{\eta}(x):=f_{j}(x)+\frac{1}{2\eta}\|x-y\|^{2}\right\},$ (11)
$\tilde{x}_{j}$ $=\mathrm{argmin}\,\left\{f_{y}^{\eta}(x):x\in\{x_{j},\tilde{x}_{j-1}\}\right\}.$ (12)
     If $f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})\leq\delta$, then stop and return $J=j,x_{J},\tilde{x}_{J}$.
  end for
@@
    $f_{j}(x)\leq f_{j+1}(x)$ and $f_{j}(x)\leq f(x)$ for $\forall x\in\mathbb{R}^{d}$ and $\forall j\geq 1$;
@@
    $f_{j}^{\eta}(x_{j})+\|x-x_{j}\|^{2}/(2\eta)\leq f_{j}^{\eta}(x)$ for $\forall x\in\mathbb{R}^{d}$ and $\forall j\geq 1$;
@@
    $\delta_{J}\leq\delta$ and $\delta_{j+1}+\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}\leq\delta_{j}$ for $\forall j\geq 1$;
@@
    $f_{y}^{\eta}(\tilde{x}_{J})-f_{y}^{\eta}(x)\leq\delta-\frac{1}{2\eta}\|x_{J}-x\|^{2}$ for $\forall x\in\mathbb{R}^{d}$;
@@
    $-\frac{1}{\eta}(x^{*}-y)\in\partial f(x^{*})$ where $\partial f$ denotes the subdifferential of $f$.
Proof a) The first inequality follows from the definition of $f_{j}$ in step 2 of Algorithm 2. The second inequality directly follows from the definition of $f_{j}$ and the convexity of $f$.
b) This statement follows from (11) and the fact that $f_{j}^{\eta}$ is $(1/\eta)$-strongly convex.
@@
$\delta_{j+1}$ $=f_{y}^{\eta}(\tilde{x}_{j+1})-f_{j+1}^{\eta}(x_{j+1})\leq f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}$
@@
This statement then follows from rearranging the terms and the definition of $f_{y}^{\eta}$ in (9)
e) This statement directly follows from the optimality condition of (9). ∎
Clearly, when Algorithm 2 terminates, the output $\tilde{x}_{J}$ is a $\delta$-solution to (9). To see this, note that, using the first inequality in Lemma 3.1(c), (11), and the fact that $f_{J}^{\eta}(\cdot)\leq f_{y}^{\eta}(\cdot)$, we have
@@
It is also easy to see that $\delta_{j}$ is computable upper bound on the gap $f_{y}^{\eta}(\tilde{x}_{j})-f_{y}^{\eta}(x^{*})$. Hence, Algorithm 2 terminates when $\delta_{j}\leq\delta$.
@@ 3.1 Complexity for Hölder smooth optimization
Proof a) It follows from the definition of $\delta_{j}$ in (13) and the definition of $\tilde{x}_{j}$ in (12) that
$\delta_{j}$ $\leq f_{y}^{\eta}(x_{j})-f_{j}^{\eta}(x_{j})=f(x_{j})-f_{j}(x_{j})$
@@
b) This statement directly follows from a) and the second inequality in Lemma 3.1(c). ∎
@@
Proof a) Using the definition of $\beta$ in (14), the assumption that $\delta_{j}>\delta$, and Lemma 3.2(b), we obtain
@@
where the last inequality is due to the definition of $j_{0}$ in (14). This contradicts with the assumption that $\delta_{j_{0}}>\delta$, and hence we prove this statement. ∎
@@
Proof Following the optimality condition of (11) with $j=1$, we have $x_{0}-x_{1}=\eta f^{\prime}(x_{0})=\eta f^{\prime}(y)$. This identity and Lemma 3.2(a) with $j=1$ then imply that the lemma holds. ∎
@@
Proof This theorem follows directly from Proposition 3.3 and Lemma 3.4. ∎
@@ 3.2 Complexity for hybrid optimization
Proof a) Following a similar argument as in the proof of Lemma 3.2(a) with (7) replaced by (8), we have
@@
and hence the statement follows. ∎
@@
Proof This lemma follows from a similar argument as in the proof of Lemma 3.4. ∎
@@
Hence, it is easy to see that $\delta_{j}\leq\delta$ if $j\geq\frac{1}{1-\tau}\log\left(\frac{2\delta_{1}}{\delta}\right)$. Using the definition of $\tau$ in (18), we have if $j$ is as in (17), then $\delta_{j}\leq\delta$. ∎
@@
Proof This theorem follows directly from Proposition 3.8 and Lemma 3.7. ∎

### v3

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex but not necessarily smooth. In sampling, a potential of the distribution $\nu(x)$ is defined as the negative log-density, which is $f(x)$ up to a constant.
@@
Optimization and sampling are two of the most important algorithmic tools at the interface of data science and computation. Optimization has been extensively studied across a wide range of fields, including machine learning, communications, and supply chain management. Over the past two decades, particular attention has been devoted to gradient-based first-order methods. Many classical ideas have been revisited and extended to large-scale optimization, such as the randomized coordinate descent method [52], the primal–dual hybrid gradient method [5], and the extragradient method [30]. Drawing samples from a given (often unnormalized) probability density plays a crucial role in many scientific and engineering problems that face uncertainty (either physically or algorithmically). Sampling algorithms are widely used in many areas such as statistical inference/estimation, operations research, physics, biology, and machine learning, etc [2, 11, 12, 16, 25, 26, 31, 64]. For instance, in Bayesian inference, one draws samples from the posterior distribution to infer its mean, covariance, or other important statistics. Sampling is also heavily used in molecular dynamics to discover new molecular structures.
@@
This work is along the recent line of research that lies in the interface of sampling and optimization [10, 62]. Indeed, sampling is closely related to optimization. On the one hand, optimization can be viewed as the limiting case of sampling from the distribution $\exp(-f(x)/T)$ as the temperature parameter $T$ (which represents the level of randomness) approaches zero. In this limit, the probability mass increasingly concentrates around the minimizers of $f(x)$. On the other hand, sampling $\nu(x)$ has an optimization interpretation [24, 67, 69]: the Langevin dynamics in space corresponds to the Fokker-Planck equation, which is the gradient flow of the relative entropy functional (with respect to $\nu$) in the space of measures with the Wasserstein metric. The popular gradient-based Markov chain Monte Carlo (MCMC) methods such as Langevin Monte Carlo (LMC) [7, 20, 56, 58], Metropolis-adjusted Langevin algorithm (MALA) [3, 57, 58], and Hamiltonian Monte Carlo (HMC) [49] resemble the gradient-based algorithms in optimization and can be viewed as the sampling counterparts of them.
@@
The goal of this paper is to develop efficient proximal algorithms to solve optimization problems (1) as well as to draw samples from potentials (2), where both $f$ in (1) and (2) lack smoothness (i.e., when $f$ does not have Lipschitz continuous gradient). In particular, we consider two settings where the convex objective/potential function $f$ is either Hölder smooth (i.e., the (sub)gradient $f^{\prime}$ is Hölder-continuous with exponent $\alpha\in[0,1]$) or a hybrid function with multiple Hölder smooth components. The core of both proximal optimization and sampling algorithms lies in the proximal map of $f$. We first develop a generic and efficient implementation of this proximal map. Building on it, we design an adaptive proximal bundle method to solve problem (1). Furthermore, by combining the proximal map of $f$ with rejection sampling, we propose a highly efficient approach to realize a proximal sampling oracle, which is used in a proximal sampling framework [33, 6] in the same spirit as the proximal point method for optimization. With those proximal oracles for optimization and sampling in hand, we are finally able to establish the complexity to sample from densities with non-smooth potentials.
@@ 2 Proximal Optimization and Sampling
Starting from the original paper [33] that proposes ASF, subsequent works have refined and extended this framework. In particular, [6] provides an improved theoretical analysis of ASF, and [70] studies Gibbs sampling based on ASF for structured log-concave distributions over networks. In Algorithm 1, sampling $y_{k}$ given $x_{k}$ in step 1 can be easily done since $\pi^{Y|X}(y\mid x_{k})={\cal N}(x_{k},\eta I)$ is a simple Gaussian distribution. Sampling $x_{k+1}$ given $y_{k}$ in step 2 is however a nontrivial task; it corresponds to the so-called restricted Gaussian oracle (RGO) for $f$ introduced in [33], which is defined as follows.
@@
A rejection sampling-based implementation of RGO for general convex nonsmooth potential function $f$ with bounded Lipschitz constant is given in [37]. If the stepsize $\eta$ is small enough, then it only takes a constant number of rejection steps to generate a sample according to RGO in expectation. Another exact realization of RGO is provided in [38] for nonconvex hybrid potential $f$ satisfying Hölder continuous conditions. It is also shown that the expected number of rejections to implement RGO is a small constant if $\eta$ is small enough. Other inexact realizations of RGO based on approximate rejection sampling are studied in [18, 13]. See Table 1 for a clear comparison. In all these implementations, a key step is realizing the proximal map (4). It is worth noting that [38] also connects ASF with other well-known Langevin-type sampling algorithms such as Langevin Monte Carlo (LMC) and Proximal Langevin Monte Carlo (PLMC) via RGO. In a nutshell, [38] shows that both LMC and PLC are instances of ASF but with approximate implementations of RGO, which always accept the sample from the proposal distribution without rejection. Hence, this provides an alternative interpretation of why the samples generated by LMC are biased, while those produced by ASF are unbiased.
Papers RGO implementation Stepsize $\eta$
[37, 38] Exact Small
[18, 13] Approximate Large
Table 1: Comparison of different RGO implementations and corresponding stepsizes.
@@
Example.
Consider the $\ell_{p}$ regression problem with data $\{(a_{i},b_{i})\}_{i=1}^{n}$ where $a_{i}\in\mathbb{R}^{d}$ and $b_{i}\in\mathbb{R}$ for $i=1,\ldots,n$,
$f(x)=\frac{1}{n}\sum_{i=1}^{n}|a_{i}^{\top}x-b_{i}|^{p},\qquad 1\leq p\leq 2.$ (9)
Define $\phi(t)=|t|^{p}$, then $\phi^{\prime}(t)=p\,\operatorname{sign}(t)\,|t|^{p-1}$ and
$f^{\prime}(x)=\frac{1}{n}\sum_{i=1}^{n}\phi^{\prime}(a_{i}^{\top}x-b_{i})\,a_{i}.$
It is shown in Lemma A.4 of Appendix A that $\phi^{\prime}$ is Hölder continuous with exponent p1p-1 and constant $p2^{2-p}$. For any $x,y\in\mathbb{R}^{d}$, let $u_{i}=a_{i}^{\top}x-b_{i}$ and $v_{i}=a_{i}^{\top}y-b_{i}$, using the Hölder continuity of $\phi^{\prime}$, we derive
$\|f^{\prime}(x)-f^{\prime}(y)\|=\Big\|\frac{1}{n}\sum_{i=1}^{n}\big(\phi^{\prime}(u_{i})-\phi^{\prime}(v_{i})\big)a_{i}\Big\|$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}|\phi^{\prime}(u_{i})-\phi^{\prime}(v_{i})|\,\|a_{i}\|\leq\frac{p\,2^{2-p}}{n}\Big(\sum_{i=1}^{n}\|a_{i}\|^{p}\Big)\|x-y\|^{p-1}.$
Hence, $f$ satisfies the Hölder smoothness condition (5) with
$\alpha=p-1,\qquad L_{\alpha}=\frac{p\,2^{2-p}}{n}\sum_{i=1}^{n}\|a_{i}\|^{p}.$
The $\ell_{p}$ regression can be extended to mixed-exponent regression as an example of the hybrid case (6), where
$f(x)=\frac{1}{n}\sum_{i=1}^{n}|a_{i}^{\top}x-b_{i}|^{p_{i}},\qquad 1\leq p_{i}\leq 2,$ (10)
and
$\alpha_{i}=p_{i}-1,\qquad L_{\alpha_{i}}=\frac{p_{i}\,2^{2-p_{i}}}{n}\|a_{i}\|^{p_{i}}.$
The above objective functions $f$ in (9) and (10) can also appear as the potential energy in Bayesian inference. Instead of minimizing $f(x)$ to obtain a point estimate (e.g., the maximum a posteriori or MAP solution), one may consider sampling $\nu(x)\propto\exp(-f(x))$ for quantifying uncertainty around the MAP solution.
[…]
@@ 3 Algorithm and Complexities for the Proximal Subproblem
Since the prox center $y$ is fixed throughout this section, we simplify the notation by writing $f_{y}^{\eta}$ as $f^{\eta}$ in this section to ease readability.
@@
0: Let $y\in\mathbb{R}^{d}$, $\eta>0$, and $\delta>0$ be given, and set $x_{0}=\tilde{x}_{0}=y$, $j=1$, and $f_{0}^{\eta}(x_{0})=$-$\infty$.
while $f^{\eta}(\tilde{x}_{j-1})-f_{j-1}^{\eta}(x_{j-1})>\delta$ do
$f_{j}(x)$ $=\max\left\{f(x_{i})+\langle f^{\prime}(x_{i}),x-x_{i}\rangle:\,0\leq i\leq j-1\right\},$ (12)
$x_{j}$ $=\mathrm{argmin}\,\left\{f_{j}^{\eta}(x):=f_{j}(x)+\frac{1}{2\eta}\|x-y\|^{2}:x\in\mathbb{R}^{d}\right\},$ (13)
$\tilde{x}_{j}$ $=\mathrm{argmin}\,\left\{f^{\eta}(x):x\in\{x_{j},\tilde{x}_{j-1}\}\right\},$ (14)
$j$ $\leftarrow j+1.$
end while
return J=j1J=j-1, $x_{J}$, and $\tilde{x}_{J}$.
@@
    $\{f_{j}\}$ serves as a sequence of non-decreasing lower approximations of $f$: $f_{j}(x)\leq f_{j+1}(x)$ and $f_{j}(x)\leq f(x)$, $\forall x\in\mathbb{R}^{d}$ and $\forall j\geq 1$;
@@
    direct consequence of (13): $f_{j}^{\eta}(x_{j})+\|x-x_{j}\|^{2}/(2\eta)\leq f_{j}^{\eta}(x)$, $\forall x\in\mathbb{R}^{d}$ and $\forall j\geq 1$;
@@
    $\{\delta_{j}\}$ is a decreasing sequence: $\delta_{J}\leq\delta$ and $\delta_{j+1}+\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}\leq\delta_{j}$, $\forall j\geq 1$;
@@
    solution guarantee for $x_{J}$ and $\tilde{x}_{J}$: $f^{\eta}(\tilde{x}_{J})-f^{\eta}(x)\leq\delta-\frac{1}{2\eta}\|x_{J}-x\|^{2}$, $\forall x\in\mathbb{R}^{d}$;
@@
    optimality condition of (11): $\frac{1}{\eta}(x^{*}-y)\in\partial f(x^{*})$ where $\partial f$ denotes the subdifferential of $f$.
Proof: a) The first inequality follows from the definition of $f_{j}$ in step 2 of Algorithm 2. The second inequality directly follows from the definition of $f_{j}$ and the convexity of $f$.
b) Noting that $f_{j}^{\eta}$ as the objective function of (13) is $(1/\eta)$-strongly convex, it thus follows from Theorem 5.25 of [1] that
$f_{j}^{\eta}(x)-f_{j}^{\eta}(x_{j})\geq\frac{1}{2\eta}\|x-x_{j}\|^{2},\quad\forall x\in\mathbb{R}^{d}.$
Hence, this statement follows.
@@
$\delta_{j+1}$ $=f^{\eta}(\tilde{x}_{j+1})-f_{j+1}^{\eta}(x_{j+1})\leq f^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}$
@@
This statement then follows from rearranging the terms and the definition of $f^{\eta}$ in (11).
e) This statement directly follows from the first-order optimality condition of (11).
Clearly, when Algorithm 2 terminates, the output $\tilde{x}_{J}$ is a $\delta$-solution to (11). To see this, note that, using the first inequality in Lemma 3.1(c), (13), and the fact that $f_{J}^{\eta}(\cdot)\leq f^{\eta}(\cdot)$, we have
@@
It is also easy to see that $\delta_{j}$ is computable upper bound on the gap $f^{\eta}(\tilde{x}_{j})-f^{\eta}(x^{*})$. Hence, Algorithm 2 terminates when $\delta_{j}\leq\delta$.
@@ 3.1 Complexity for Hölder Smooth Optimization
Proof: a) It follows from the definition of $\delta_{j}$ in (15) and the definition of $\tilde{x}_{j}$ in (14) that
[…]
@@
b) This statement directly follows from a) and the second inequality in Lemma 3.1(c).
@@
Proof: a) Using the definition of $\beta$ in (16), the assumption that $\delta_{j}>\delta$, and Lemma 3.2(b), we obtain
@@
where the last inequality is due to the definition of $j_{0}$ in (16). This contradicts with the assumption that $\delta_{j_{0}}>\delta$, and hence we prove this statement.
@@
Proof: Following the optimality condition of (13) with $j=1$, we have $x_{0}-x_{1}=\eta f^{\prime}(x_{0})=\eta f^{\prime}(y)$. This identity and Lemma 3.2(a) with $j=1$ then imply that the lemma holds.
@@
Proof: This theorem follows directly from Proposition 3.3 and Lemma 3.4.
@@ 3.2 Complexity for Hybrid Optimization
Proof: a) Following a similar argument as in the proof of Lemma 3.2(a) with (7) replaced by (8), we have
@@
and hence the statement follows.
@@
Proof: This lemma follows from a similar argument as in the proof of Lemma 3.4.
@@
Hence, it is easy to see that $\delta_{j}\leq\delta$ if $j\geq\frac{1}{1-\tau}\log\left(\frac{2\delta_{1}}{\delta}\right)$. Using the definition of $\tau$ in (20), we have if $j$ is as in (19), then $\delta_{j}\leq\delta$.
@@
Proof: This theorem follows directly from Proposition 3.8 and Lemma 3.7.
3.3 Implementation of Algorithm 2
This subsection presents the simulation results of Algorithm 2 on solving the regularized subproblem (11) for two objective functions $f$: quadratic programming (QP) and $\ell_{p}$ regression. In both cases, the subgradient $f^{\prime}$ is computed by automatic differentiation via Zygote.jl [22], and the subproblem (13) is reformulated as a QP and solved using Clarabel.jl [19]. Numerical simulations are conducted on an i9-13900k desktop with 64 GB of RAM
Quadratic Programming We first consider the unconstrained QP problem
$f(x)=\frac{1}{2}x^{\top}Qx+\langle c,x\rangle$
where $Q\in\mathbb{S}^{d}_{+}$ and $c\in\mathbb{R}^{d}$. We generate $Q=AA^{\top}/\|AA^{\top}\|_{\infty}$ where $A\in\mathbb{R}^{d\times d}$ has normally distributed entries and $\|AA^{\top}\|_{\infty}=\max_{ij}|(AA^{\top})_{ij}|$ is the entrywise infinity norm. The linear term $c$ and point $y$ are also entrywise normally distributed. The dimension $d$ is set to be $1000$.
Note that (11) for QP has the closed-form solution
$\underset{x\in\mathbb{R}^{n}}{\mathrm{argmin}\,}\left\{\frac{1}{2}x^{\top}Qx+\langle c,x\rangle+\frac{1}{2\eta}\|x-y\|^{2}\right\}=(Q+\eta^{-1}I)^{-1}(\eta^{-1}y-c),$
hence we can compare the progress of Algorithm 2 against the true minimum. We run Algorithm 2 until the condition $\delta_{j}<10^{-6}$ is satisfied. Fig. 1 shows the function value decrease of the minimum value iterate $f_{y}^{\eta}(\tilde{x}_{j})$ versus the optimal value $f_{y}^{\eta}(x^{*})$ with varying $\eta$. Noting that Algorithm 2 requires more iterations as $\eta$ increases, this observation is consistent with the complexity bound (proportional to $\eta$) stated in Theorem 3.5.
Figure 1: Proximal subproblem progress of Algorithm 2 in quadratic programming.
$\ell_{p}$ Regression We next consider $\ell_{p}$ regression, where the objective $f$ is of the form
$f(x)=\|Ax-b\|_{p}^{p},$
where $A\in\mathbb{R}^{n\times d}$ and $b\in\mathbb{R}^{n}$ have normally distributed entries and $A$ is again divided by its entrywise infinity norm. We set $d=100$ and $n=500$ for testing. The point $y\in\mathbb{R}^{d}$ is entrywise normally distributed, and is identical for all $p$ values tested. Algorithm 2 is terminated when $\delta_{j}<10^{-6}$. Fixing $\eta=1.0$, Fig. 2 shows the trajectory of the gap $\delta_{j}$ for varying $p$ values.
Figure 2: Proximal subproblem progress of Algorithm 2 in $\ell_{p}$ regression.